# An Equivalence Result on the Order of Differentiability in Frobenius’ Theorem

### v1

@@ 1 Introduction
This paper aims to solve this asymmetric result. Even in the above example, the integral curve must be $C^{2}$. [^3] A level set of a solution to (1) is sometimes called an integral manifold. We focus on this feature, and show that for a locally Lipschitz function $g:U\to\mathbb{R}^{n}\setminus\{0\}$ that satisfies (2) almost everywhere, there exists a solution to (1) such that any level set of $u$ is an n1n-1 dimensional $C^{1}$ manifold, and $g$ is $C^{k}$ if and only if for any $x^{*}\in U$, there is a solution $(u,\lambda)$ around $x^{*}$ such that $u$ is $C^{k}$ and any level set of $u$ is an n1n-1-dimensional $C^{k+1}$ manifold.
@@ 2.3 Differentiable Manifolds
Recall the definition of differentiable manifolds. We use only regular submanifolds of some Euclidean space, and thus we can use the definition of manifold that appears in Milnor [13] or Guillemin and Pollack [6]. [^5] The proofs of facts treated in this subsection are found in Chapter 1 of Guillemin and Pollack [6]. First, suppose that $X\subset\mathbb{R}^{N}$ and $f:X\to\mathbb{R}^{M}$. We say that $f$ is $C^{k}$ if for any $x\in X$, there exists $r>0$ and a $C^{k}$ mapping $F:B_{r}(x)\to\mathbb{R}^{M}$ such that if $y\in U\cap B_{r}(x)$, then $F(y)=f(y)$. A bijection $f:X\to Y$ is called a $C^{k}$ diffeomorphism if $f$ and $f^{-1}$ are $C^{k}$, and two sets $X,Y$ are said to be $C^{k}$ diffeomorphic if there is a $C^{k}$ diffeomorphism $f:X\to Y$. A set $X\subset\mathbb{R}^{N}$ is called an $n$-dimensional $C^{k}$ manifold if, for any $x\in X$, there exists an open neighborhood $V\subset X$ of $x$ that is $C^{k}$ diffeomorphic to some open set $U$ in $\mathbb{R}^{n}$. The corresponding $C^{k}$ diffeomorphism $\phi:U\to V$ is called a local parametrization around $x$, and the inverse $\phi^{-1}$ is called a local coordinate. The dimension $n$ of $X$ is denoted by $\dim X$.
@@
Let $X,Y$ be $C^{1}$ manifolds, and $f:X\to Y$ be $C^{1}$. A point $y\in Y$ is called a regular value if for each $x\in f^{-1}(y)$, $df_{x}$ is onto. It is well-known that if $y$ is a regular value and $f^{-1}(y)$ is nonempty, then $f^{-1}(y)$ is a $C^{1}$ manifold whose dimension is $\dim X-\dim Y$. This result is called the preimage theorem.
@@ 2.4 Ordinary Differential Equations
where $U\subset\mathbb{R}\times\mathbb{R}^{N}\times\mathbb{R}^{M}$ is open and $h:U\to\mathbb{R}^{N}$ is continuous. A continuously differentiable function $x(t)$ is called a solution to (3) if it is defined on an interval $I\subset\mathbb{R}$ that includes $t_{0}$, [^6] We call a set $I\subset\mathbb{R}$ an interval if $I$ is a convex subset of $\mathbb{R}$ that includes at least two different points. $\dot{x}(t)=h(t,x(t),z)$ for all $t\in I$, and $x(t_{0})=x_{0}$. If there are two solutions $x(t),y(t)$ to (3), $x(t)$ is called an extension of $y(t)$ if the domain of $x(t)$ includes that of $y(t)$ and $x(t)=y(t)$ if both are defined at $t$. A solution $x(t)$ to (3) is called a nonextendable solution if there is no extension of $x(t)$ except $x(t)$ itself. It is known that if $h$ is locally Lipschitz, then for all $(t_{0},x_{0},z)\in U$, there uniquely exists a nonextendable solution $x(t;z,x_{0})$ to (3). The mapping $(t;z,x_{0})\mapsto x(t;z,x_{0})$ is called the solution function of (3). It is also well-known that the domain of the solution function is open, and the solution function is continuous on its domain. Moreover, if $h$ is locally Lipschitz, then $c$ is also locally Lipschitz, and if $h$ is $C^{k}$, then $c$ is also $C^{k}$. [^7] Almost all assertions in this paragraph are proved in famous textbooks. See, for example, Pontryagin [15], Hartman [7], or Coddington and Levinson [2]. Only the local Lipschitz property of $c$ is a little hard to find a proof, but it appears in Subsection 7.1 of Hosoya [10].
@@ 2.5 Convex Analysis
A function $f:C\to\mathbb{R}$ is said to be concave if f-f is convex. In other words, $f$ is concave if and only if for any $x,y\in C$ and $t\in[0,1]$,
@@
Sometimes, a relaxed condition for convexity is useful. Suppose that $C\subset\mathbb{R}^{n}$ is convex, and let $f:C\to\mathbb{R}$. Then, $f$ is said to be quasi-convex if for any $x\in C$, the lower contour set $L(x)=\{y\in C|f(y)\leq f(x)\}$ is convex. This is equivalent to the following condition: for any $x,y\in C$ and $t\in[0,1]$,
@@
$f$ is said to be quasi-concave if f-f is quasi-convex. In other words, $f$ is quasi-concave if and only if for any $x,y\in C$ and $t\in[0,1]$,
@@
Suppose that $U$ is open and convex, and $f:U\to\mathbb{R}$ is convex. For each $x\in U$, define
@@ 3.1 Normal Solutions to the Total Differential Equation
Proof.Suppose that $V\subset U$ is open and a pair of functions $u:V\to\mathbb{R},\lambda:V\to\mathbb{R}$ satisfy all requirements in this theorem, and choose any $x\in V$. Define $X=u^{-1}(u(x))$. Because $u$ is nondegenerate, $u(x)$ is a regular value of $u$. By the preimage theorem, $X$ is an n1n-1 dimensional $C^{1}$ manifold, and thus $u$ is a normal solution. Let $v\in T_{x}(X)$, and choose a $C^{1}$ function $c:(-\varepsilon,\varepsilon)\to X$ such that $c(0)=x$ and $\dot{c}(0)=v$. Then, $u(c(t))\equiv u(x)$, and thus
@@
Conversely, suppose that $u:V\to\mathbb{R}$ is a $C^{1}$ normal solution to (1). Then, $\nabla u(x)\neq 0$ everywhere. As in the last paragraph, we can show that for $X=u^{-1}(u(x))$, $g(x)$ is orthogonal to $T_{x}(X)$. Because both $g(x)$ and $\nabla u(x)$ are orthogonal to the n1n-1 dimensional space $T_{x}(X)$, $\nabla u(x)$ is proportional to $g(x)$, and thus there exists $\lambda(x)\in\mathbb{R}$ such that $\nabla u(x)=\lambda(x)g(x)$. Choose any $i$ such that $g_{i}(x)\neq 0$. Then,
@@ 3.2 Key Lemmas
Suppose that $U\subset\mathbb{R}^{n}$ is open and $g:U\to\mathbb{R}^{n}\setminus\{0\}$ is locally Lipschitz. Moreover, suppose that $u:V\to\mathbb{R}$ is a normal solution to (1). Choose any $x\in V$ and $v\in\mathbb{R}^{n}$ such that $g(x)\cdot v\neq 0$, and define $c_{x}(t)=u(x+tv)$. Then, the following results hold.
@@
Fourth, suppose that $V$ is connected and $g(x)\cdot v\neq 0$ for all $x\in U$. Let $U_{1}$ be the set of all $x\in U$ such that $c_{x}(t)$ is increasing around 0. By the above arguments, we have that both $U_{1}$ and $V\setminus U_{1}$ are open. Because $V$ is connected, either $U_{1}=V$ or $U_{1}=\emptyset$, and (4) holds.
@@
We say that $g$ satisfies Jacobi’s integrability condition if for any $i,j,k\in\{1,...,n\}$, (6) holds almost everywhere.
@@
$\frac{\partial E^{x_{n}}}{\partial x_{i}}(\tilde{y})=f_{i}(\tilde{y},E^{x_{n}}(\tilde{y})),\ E^{x_{n}}(\tilde{x})=x_{n}.$ (7)
@@
Choose any $\tilde{y}\in W$ and $x_{n}\in I$, and let $)y=(\tilde{y},E^{x_{n}}(\tilde{y}$)). Because $c_{i}(t)=(\tilde{y}+t\tilde{e}_{i},E^{x_{n}}(\tilde{y}+t\tilde{e}_{i}))$ is a $C^{1}$ curve on $X$, we have that $\dot{c}_{i}(0)\in T_{y}(X)$, and thus $\dot{c}_{i}(0)\cdot g(y)=0$. Hence,
$g_{i}(y)+g_{n}(y)\frac{\partial E^{x_{n}}}{\partial x_{i}}(\tilde{y})=0,$
@@ 3.3 Invariance of the Level Sets
Suppose that $I\subset\mathbb{R}$ is an open interval, and let $U=I^{n}$ and $g:U\to\mathbb{R}^{n}_{+}\setminus\{0\}$ is locally Lipschitz. Then, any regular solution $u$ to (1) defined on $U$ is increasing, and any level set of $u$ is connected. In particular, if $u_{1}:U\to\mathbb{R}$ and $u_{2}:U\to\mathbb{R}$ are normal solutions to (1). Then, $u_{2}$ is a monotone transform of $u_{1}$.
@@
Second, choose $x,y\in U$ such that $x\neq y$ and $u(x)=u(y)$, and let $X=u^{-1}(u(x))$. Let $z_{i}=\min\{x_{i},y_{i}\}$ and $w_{i}=\max\{x_{i},y_{i}\}$. Define
@@
and v=wzv=w-z. Then, $v\gneq 0$, and because $u$ is strictly increasing and continuous, for each $t\in[0,2]$, there uniquely exists $s(t)\in[0,1]$ such that $u(x(t)+s(t)v)=u(x)$. By Berge’s theorem, we can show that $s(t)$ is continuous, and thus $c(t)=x(t)+s(t)v$ is a continuous function from $[0,2]$ into $X$ such that $c(0)=x$ and $c(2)=y$, which implies that $X$ is path connected, and thus it is connected.
@@ 3.4 Main Result
Suppose that $U\subset\mathbb{R}^{n}$ is open and $g:U\to\mathbb{R}^{n}\setminus\{0\}$ is nonempty. Then, $g$ satisfies Jacobi’s integrability condition if and only if, for any $x^{*}\in U$, there exists a normal solution $u:V\to\mathbb{R}$ to (1) around $x^{*}$. Moreover, if $g$ is $C^{k}$ and satisfies Jacobi’s integrability condition, then for any $x^{*}\in U$, there exists a $C^{k}$ normal solution $u:V\to\mathbb{R}$ to (1) around $x^{*}$ such that any level set of $u$ is an n1n-1 dimensional $C^{k+1}$ manifold.
@@
Let $c(t;\tilde{x},z)$ be the solution function of (9). It is clear that $c(t;\tilde{x}^{*},x_{n}^{*})\equiv x_{n}^{*}$, which is defined on $\mathbb{R}$. Because the domain $U^{*}$ of the solution function is open and includes $0,1]\times\{x^{*}\}$, there exists $\varepsilon>0$ such that $0,1]\times\prod_{i=1}^{n}[x_{i}^{*}-\varepsilon,x_{i}^{*}+\varepsilon]\subset U^{*}$. Define
@@
Next, we show that $u$ is Lipschitz on $V$. To show this, we first show that $u$ is Lipschitz in $x_{n}$ on $V$. Suppose that $\tilde{x}\in\prod_{i=1}^{n-1}(x_{i}^{*}-\delta,x_{i}^{*}+\delta)$, and let $\tilde{x}(t)=(1-t)\tilde{x}^{*}+t\tilde{x}$. If $z,z^{\prime}\in(x_{n}^{*}-\varepsilon,x_{n}^{*}+\varepsilon)$ and $z^{\prime}>z$, then $(\tilde{x}(t),c(t;\tilde{x},z)),(\tilde{x}(t),c(t;\tilde{x},z^{\prime}))\in\prod_{i=1}^{n-1}[x_{i}^{*}-\varepsilon,x_{i}^{*}+\varepsilon]\times C$ for any $t\in[0,1]$. Moreover, $c(t;\tilde{x},z^{\prime})>c(t;\tilde{x},z)$ for all $t\in[0,1]$. Thus,
@@
To summarize these results, we obtain a normal solution $u:V\to\mathbb{R}$ to (1) around $x^{*}$. Therefore, the first assertion of this theorem holds. Now, suppose that $g$ is $C^{k}$. Then, $f$ is also $C^{k}$. By the bootstrap argument, we can show that $E^{z}$ is $C^{k+1}$. Therefore, for any $z$, $u^{-1}(z)$ is an n1n-1 dimensional $C^{k+1}$ manifold. Choose $x\in V$, and let $z=u(x)$. Then,
$c(1;\tilde{x},z)=x_{n}.$

### v2

@@ 1 Introduction
This paper aims to solve this asymmetric result. Even in the above example, the integral curve must be $C^{2}$. [^3] A level set of a solution to (1) is sometimes called an integral manifold. We focus on this feature, and show that, 1) for a locally Lipschitz function $g:U\to\mathbb{R}^{n}\setminus\{0\}$, it satisfies (2) almost everywhere if and only if, for any $x^{*}\in U$, there exists a solution $(u,\lambda)$ to (1) around $x^{*}$ such that $u$ is locally Lipschitz and any level set of $u$ is an n1n-1 dimensional $C^{1}$ manifold, and 2) if $g$ is $C^{k}$, then $g$ satisfies (2) everywhere if and only if, for any $x^{*}\in U$, there is a solution $(u,\lambda)$ around $x^{*}$ such that $u$ is $C^{k}$ and any level set of $u$ is an n1n-1-dimensional $C^{k+1}$ manifold.
@@ 2.3 Differentiable Manifolds
Recall the definition of differentiable manifolds. We use only regular submanifolds of some Euclidean space, and thus we can use the definition of manifold that appears in Milnor [13] or Guillemin and Pollack [6]. [^5] The proofs of facts treated in this subsection are found in Chapter 1 of Guillemin and Pollack [6]. First, suppose that $X\subset\mathbb{R}^{N}$ and $f:X\to\mathbb{R}^{M}$. We say that $f$ is $C^{k}$ if, for any $x\in X$, there exists $r>0$ and a $C^{k}$ mapping $F:B_{r}(x)\to\mathbb{R}^{M}$ such that if $y\in U\cap B_{r}(x)$, then $F(y)=f(y)$. A bijection $f:X\to Y$ is called a $C^{k}$ diffeomorphism if $f$ and $f^{-1}$ are $C^{k}$, and two sets $X,Y$ are said to be $C^{k}$ diffeomorphic if there is a $C^{k}$ diffeomorphism $f:X\to Y$. A set $X\subset\mathbb{R}^{N}$ is called an $n$-dimensional $C^{k}$ manifold if, for any $x\in X$, there exists an open neighborhood $V\subset X$ of $x$ that is $C^{k}$ diffeomorphic to some open set $U$ in $\mathbb{R}^{n}$. The corresponding $C^{k}$ diffeomorphism $\phi:U\to V$ is called a local parametrization around $x$, and the inverse $\phi^{-1}$ is called a local coordinate. The dimension $n$ of $X$ is denoted by $\dim X$.
@@
Let $X,Y$ be $C^{1}$ manifolds, and $f:X\to Y$ be $C^{1}$. A point $y\in Y$ is called a regular value if, for each $x\in f^{-1}(y)$, $df_{x}$ is onto. It is well-known that if $y$ is a regular value and $f^{-1}(y)$ is nonempty, then $f^{-1}(y)$ is a $C^{1}$ manifold whose dimension is $\dim X-\dim Y$. This result is called the preimage theorem.
@@ 2.4 Ordinary Differential Equations
where $U\subset\mathbb{R}\times\mathbb{R}^{N}\times\mathbb{R}^{M}$ is open and $h:U\to\mathbb{R}^{N}$ is continuous. A continuously differentiable function $x(t)$ is called a solution to (3) if it is defined on an interval $I\subset\mathbb{R}$ that includes $t_{0}$, [^6] We call a set $I\subset\mathbb{R}$ an interval if $I$ is a convex subset of $\mathbb{R}$ that includes at least two different points. $\dot{x}(t)=h(t,x(t),z)$ for all $t\in I$, and $x(t_{0})=x_{0}$. If there are two solutions $x(t),y(t)$ to (3), $x(t)$ is called an extension of $y(t)$ if the domain of $x(t)$ includes that of $y(t)$ and $x(t)=y(t)$ if both are defined at $t$. A solution $x(t)$ to (3) is called a nonextendable solution if there is no extension of $x(t)$ except $x(t)$ itself. It is known that if $h$ is locally Lipschitz, then for all $(t_{0},x_{0},z)\in U$, there uniquely exists a nonextendable solution $x(t;z,x_{0})$ to (3). The mapping $(t;z,x_{0})\mapsto x(t;z,x_{0})$ is called the solution function of (3). It is also well-known that if $h$ is locally Lipschitz, then the domain of the solution function is open, and the solution function is locally Lipschitz. Moreover, if $h$ is $C^{k}$, then $x$ is also $C^{k}$. [^7] Almost all assertions in this paragraph are proved in famous textbooks. See, for example, Pontryagin [15], Hartman [7], or Coddington and Levinson [2]. Only the local Lipschitz property of $x$ is a little hard to find a proof, but it appears in Subsection 7.1 of Hosoya [10].
@@ 2.5 Convex Analysis
A function $f:C\to\mathbb{R}$ is said to be concave if f-f is convex. In other words, $f$ is concave if and only if, for any $x,y\in C$ and $t\in[0,1]$,
@@
Sometimes, a relaxed condition for convexity is useful. Suppose that $C\subset\mathbb{R}^{n}$ is convex, and let $f:C\to\mathbb{R}$. Then, $f$ is said to be quasi-convex if, for any $x\in C$, the lower contour set $L(x)=\{y\in C|f(y)\leq f(x)\}$ is convex. This is equivalent to the following condition: for any $x,y\in C$ and $t\in[0,1]$,
@@
$f$ is said to be quasi-concave if f-f is quasi-convex. In other words, $f$ is quasi-concave if and only if, for any $x,y\in C$ and $t\in[0,1]$,
@@
Suppose that $U\subset\mathbb{R}^{n}$ is open and convex, and $f:U\to\mathbb{R}$ is convex. For each $x\in U$, define
@@ 3.1 Normal Solutions to the Total Differential Equation
Proof.Suppose that $V\subset U$ is open and a pair of functions $u:V\to\mathbb{R},\lambda:V\to\mathbb{R}$ satisfy all requirements in this theorem, and choose any $x\in V$. Define $X=u^{-1}(u(x))$. Because $u$ is nondegenerate, $u(x)$ is a regular value of $u$. By the preimage theorem, $X$ is an n1n-1 dimensional $C^{1}$ manifold. Let $v\in T_{x}(X)$, and choose a $C^{1}$ function $c:(-\varepsilon,\varepsilon)\to X$ such that $c(0)=x$ and $\dot{c}(0)=v$. Then, $u(c(t))\equiv u(x)$, and thus
@@
Conversely, suppose that $u:V\to\mathbb{R}$ is a $C^{1}$ normal solution to (1). Then, $\nabla u(x)\neq 0$ everywhere. As in the last paragraph, we can show that for $X=u^{-1}(u(x))$, $g(x)$ is orthogonal to $T_{x}(X)$. Because both $g(x)$ and $\nabla u(x)$ are orthogonal to the n1n-1 dimensional space $T_{x}(X)$, $\nabla u(x)$ is proportional to $g(x)$, and thus there exists $\lambda(x)\in\mathbb{R}$ such that $\nabla u(x)=\lambda(x)g(x)$. Choose any $i$ such that $g_{i}(x)\neq 0$ and $\frac{\partial u}{\partial x_{i}}(x)\neq 0$. Then,
@@ 3.2 Key Lemmas
Suppose that $U\subset\mathbb{R}^{n}$ is open and $g:U\to\mathbb{R}^{n}\setminus\{0\}$ is locally Lipschitz. Moreover, suppose that $u:V\to\mathbb{R}$ is a normal solution to (1)(1). Choose any $x\in V$ and $v\in\mathbb{R}^{n}$ such that $g(x)\cdot v\neq 0$, and define $c_{x}(t)=u(x+tv)$. Then, the following results hold.
@@
Fourth, suppose that $V$ is connected and $g(x)\cdot v\neq 0$ for all $x\in U$. Let $U_{1}$ be the set of all $x\in V$ such that $c_{x}(t)$ is increasing around 0. By the above arguments, we have that both $U_{1}$ and $V\setminus U_{1}$ are open. Because $V$ is connected, either $U_{1}=V$ or $U_{1}=\emptyset$, and (4) holds.
@@
We say that $g$ satisfies Jacobi’s integrability condition if, for any $i,j,k\in\{1,...,n\}$, (6) holds almost everywhere.
@@
$\frac{\partial E^{x_{n}}}{\partial x_{i}}(\tilde{x})=f_{i}(\tilde{x},E^{x_{n}}(\tilde{x})),\ E^{x_{n}}(\tilde{x}^{*})=x_{n}.$ (7)
@@
Choose any $\tilde{x}\in W$ and $x_{n}\in I$, and let $)z=(\tilde{x},E^{x_{n}}(\tilde{x}$)). Because $c_{i}(t)=(\tilde{x}+t\tilde{e}_{i},E^{x_{n}}(\tilde{x}+t\tilde{e}_{i}))$ is a $C^{1}$ curve on $X$, we have that $\dot{c}_{i}(0)\in T_{y}(X)$, and thus $\dot{c}_{i}(0)\cdot g(y)=0$. Hence,
$g_{i}(z)+g_{n}(z)\frac{\partial E^{x_{n}}}{\partial x_{i}}(\tilde{x})=0,$
@@ 3.3 Invariance of the Level Sets
Suppose that $I\subset\mathbb{R}$ is an open interval, and let $U=I^{n}$ and $g:U\to\mathbb{R}^{n}_{+}\setminus\{0\}$ is locally Lipschitz. Then, any regular solution $u$ to (1) defined on $U$ is increasing, and any level set of $u$ is connected. In particular, if $u_{1}:U\to\mathbb{R}$ and $u_{2}:U\to\mathbb{R}$ are normal solutions to (1), then $u_{2}$ is a monotone transform of $u_{1}$.
@@
Second, choose $x,y\in U$ such that $x\neq y$ and $u(x)=u(y)$, and let $X=u^{-1}(u(x))$. Let $z_{i}=\min\{x_{i},y_{i}\}$ and $w_{i}=\max\{x_{i},y_{i}\}+1$. Define
@@
and v=wzv=w-z. Then, $v\gg 0$, and because $u$ is increasing and continuous, for each $t\in[0,2]$, there uniquely exists $s(t)\in[0,1]$ such that $u(x(t)+s(t)v)=u(x)$. By Berge’s theorem, we can show that $s(t)$ is continuous, and thus $c(t)=x(t)+s(t)v$ is a continuous function from $[0,2]$ into $X$ such that $c(0)=x$ and $c(2)=y$, which implies that $X$ is path connected, and thus it is connected.
@@ 3.4 Main Result
Suppose that $U\subset\mathbb{R}^{n}$ is open and $g:U\to\mathbb{R}^{n}\setminus\{0\}$ is locally Lipschitz. Then, $g$ satisfies Jacobi’s integrability condition if and only if, for any $x^{*}\in U$, there exists a normal solution $u:V\to\mathbb{R}$ to (1) around $x^{*}$. Moreover, when $g$ is $C^{k}$, then it satisfies Jacobi’s integrability condition if and only if, for any $x^{*}\in U$, there exists a $C^{k}$ normal solution $u:V\to\mathbb{R}$ to (1) around $x^{*}$ such that any level set of $u$ is an n1n-1 dimensional $C^{k+1}$ manifold.
@@
Let $c(t;\tilde{x},z)$ be the solution function of (9). It is clear that $c(t;\tilde{x}^{*},x_{n}^{*})\equiv x_{n}^{*}$, which is defined on $\mathbb{R}$. Because the domain $U^{*}$ of the solution function $c$ is open and includes $0,1]\times\{x^{*}\}$, there exists $\varepsilon>0$ such that $0,1]\times\prod_{i=1}^{n}[x_{i}^{*}-\varepsilon,x_{i}^{*}+\varepsilon]\subset U^{*}$. Define
@@
Next, we show that $u$ is Lipschitz on $V$. To show this, we first show that $u$ is Lipschitz in $x_{n}$ on $V$. Suppose that $\tilde{x}\in\prod_{i=1}^{n-1}[x_{i}^{*}-\delta,x_{i}^{*}+\delta]$, and let $\tilde{x}(t)=(1-t)\tilde{x}^{*}+t\tilde{x}$. If $z,z^{\prime}\in[x_{n}^{*}-\varepsilon,x_{n}^{*}+\varepsilon]$ and $z^{\prime}>z$, then $(\tilde{x}(t),c(t;\tilde{x},z)),(\tilde{x}(t),c(t;\tilde{x},z^{\prime}))\in\prod_{i=1}^{n-1}[x_{i}^{*}-\varepsilon,x_{i}^{*}+\varepsilon]\times C$ for any $t\in[0,1]$. Moreover, $c(t;\tilde{x},z^{\prime})>c(t;\tilde{x},z)$ for all $t\in[0,1]$. Thus,
@@
To summarize these results, we obtain a normal solution $u:V\to\mathbb{R}$ to (1) around $x^{*}$. Therefore, the first assertion of this theorem holds. Now, suppose that $g$ is $C^{k}$. Then, $f$ is also $C^{k}$. By the bootstrap argument, we can show that $E^{z}$ is $C^{k+1}$. Therefore, for any $z$, $u^{-1}(z)$ is an n1n-1 dimensional $C^{k+1}$ manifold. Choose $x\in V$. Then, for each $z\in(x_{n}^{*}-\varepsilon,x_{n}^{*}+\varepsilon)$,
$u(x)=z\Leftrightarrow c(1;\tilde{x},z)=x_{n}.$